OpenStudy (anonymous):

find the derivative of cos2theta

OpenStudy (anonymous):

can you please show the step-by-step process please? thankyou

OpenStudy (anikhalder):

First, tell me what shall be the derivative of cosx?

hartnn (hartnn):

and you'll need chain rule, which i guess you already know

OpenStudy (anonymous):

\[(-2)\sin (2\theta)\]

OpenStudy (anonymous):

uhm, my answer for this is -4sintheta

OpenStudy (anonymous):

yeah @nitz :)

OpenStudy (anonymous):

Hi again @hartnn :)

hartnn (hartnn):

hello :)

hartnn (hartnn):

how did u get -4 ?

OpenStudy (anonymous):

2 * -2 ?

OpenStudy (anonymous):

here's the real question. \[\frac{ d }{ d \Theta } (\sin \Theta * \cos2\Theta) =\]

hartnn (hartnn):

-2, yes where does 2 come from ?

hartnn (hartnn):

ohh

OpenStudy (anonymous):

uhm, -sin2\[-\sin2\Theta * 2\] <<< chain rule?

OpenStudy (anonymous):

that's a composite function, right? O_O

hartnn (hartnn):

yes so, derivative of cos 2theta = - sin2theta * derivative of 2 theta = -2 sin 2 theta is correct :)

hartnn (hartnn):

and whats your final answer for the actual question u posted?

OpenStudy (anonymous):

im not sure actually. \[\cos \Theta*\cos2\Theta-4\sin \Theta*\sin \Theta\]

hartnn (hartnn):

once again, where does that -4 come from :P

OpenStudy (anonymous):

i keep on checking it on wolframalpha.com but i don't get the same answer :(

hartnn (hartnn):

your first term is correct

OpenStudy (anonymous):

uhm, come again please? where does my 4 come from? O_O

hartnn (hartnn):

2nd term totally incorrect

OpenStudy (anonymous):

im actually using product rule here. what should i use then?

hartnn (hartnn):

for 2nd term you'll keep sin theta as it as and derivate os 2theta

hartnn (hartnn):

right ?

hartnn (hartnn):

i meant derivate cos 2theta

hartnn (hartnn):

\(\large \cos \theta \cos 2\theta + \sin \theta \dfrac{d}{dx}\cos 2\theta\) you get this ?

hartnn (hartnn):

i meant d/dtheta -_-

OpenStudy (anonymous):

ohh. yeah, i get it

hartnn (hartnn):

good, now we solved derivative of cos 2theta = -2 sin 2 theta just plug this in

hartnn (hartnn):

\(\large \cos \theta \cos 2\theta + \sin \theta (-2\sin 2\theta)\) simplify

OpenStudy (anonymous):

oh. okay.

OpenStudy (anonymous):

we can further simplify that?

OpenStudy (anonymous):

but this, is something wrong here? http://www.wolframalpha.com/input/?i=%28sin+theta+*+cos+%282theta%29%29+find+derivative

hartnn (hartnn):

nothing much to simplify cos t cos 2t -2 sin t sin 2t wolf gave answer in terms of 3theta, do you require answer in terms of 3theta ?

hartnn (hartnn):

else, cos t cos 2t -2 sin t sin 2t can be totally acceptable final answer

OpenStudy (anonymous):

oh. okay. thankyou :)

OpenStudy (anonymous):

and can u pls help me with evaluating derivatives? :)

OpenStudy (anikhalder):

today i was going through this book..calculus for the clueless by bob millers part 1...perfect fit for me..you might try it out..not saying that you are clueless..but it's helpful

OpenStudy (anonymous):

Hahaha. Sure thing x thanks

OpenStudy (anonymous):

here are some questions @hartnn :)

OpenStudy (anonymous):

\[2. g(s) = \sqrt[3]{s} + s^{2}; s'(3)=\]

OpenStudy (anonymous):

\[3. h(t)=\sqrt{3t ^{2}-5t}; h' (\frac{ 1 }{ 6 }) = \]

OpenStudy (anonymous):

looks like you should just find the respected derivatives and plug 3 on the first and 1/6 on the second into them

OpenStudy (anonymous):

ohhh.

OpenStudy (anonymous):

is that a typo on the first one? the function is g(s) but the second part is s'(3)

OpenStudy (anonymous):

oh, i think so. but my worksheet says its 3. lol its probably a typo ;) sorry

OpenStudy (anonymous):

the 3 doestn bother me, its the fact that its wanting the derivative of the variable term. maybe someone else could make better sense of it than me. the second one is completely straightforward though

OpenStudy (anonymous):

\[a. \frac{ d }{ dx }| _{x=1} (\sqrt{x ^{3}-1}) =\] the subscript is x= 1

OpenStudy (anonymous):

is this a new question?

OpenStudy (anonymous):

\[b. \frac{ d }{ dx }| _{x=\frac{ 1 }{ 6 }} (3x ^{2}-4x+2) =\] subscript is x=1/6

OpenStudy (anonymous):

@lonnie455rich yep :)

OpenStudy (anonymous):

\[c. \frac{ d }{ dr }| _{r=2} [(5r \times \cos \sqrt{r})]=\] subscript is r=2

OpenStudy (anonymous):

your subscripts look like your restrictions on x. I guess I don't understand the question? are you just trying to find the derivatives?

OpenStudy (anonymous):

uhm, it says here 'evaluate the derivatives at given points'

OpenStudy (anonymous):

d. g(s)=s√3+s2; g′(3)=

OpenStudy (anonymous):

e. g(s)=s√(3t^2 - 5t); h/(1/6)=

OpenStudy (anonymous):

ahhh use the limit definition (f(x+h)-f(x))/h

OpenStudy (anonymous):

you may know h as delta x if that makes more sense to you

OpenStudy (anonymous):

uhm, okay. imma try understand it :)

OpenStudy (anonymous):

uhm, i dont get it. what will i use to substitute and stuff? :(

OpenStudy (anonymous):

that should have been your first chapter in learning derivatives, no? do b first since its easiest to check. then once you get the hand of it. the rest aren't so bad. here is an attachment of when I was learning this a few months ago

OpenStudy (anonymous):

sorry, i've seen it in my notes here and im trying to remember it. lol

hartnn (hartnn):

if its not mentioned to use limit definition of derivative, then we can use standard formulas

hartnn (hartnn):

\( \large g(s) =\sqrt[3]s +s^2 = s^{1/3}+s^2\) now we use the x^n formula of derivative, i think you know it, so i'll directly write next step \(\large g'(s) = (1/3)s^{-2/3}+2s\) to get g'(3), just plug in s= 3 in that and simplify :)

hartnn (hartnn):

you wanted me to solve all 7 of them :O

OpenStudy (anonymous):

ahh so if you have a point. for x you don't have to use the limit definition? just a nother form of f'(3) of f(x)=2x^2 I solved a lot of problems the hard way lol.

hartnn (hartnn):

haha, as i said, use definition only if specifically mentioned :)

OpenStudy (anonymous):

see, I just thought that if you to evaluate at a point you had to. its funny my intuition when I saw f'(3) was to compute the derivative and plug it in but when he said at a point I thought limit definition. frustrating sometimes. not knowing shortcuts.lol